# Relating Checkpoint Update Probabilities to Momentum Parameters in Single-Loop Variance Reduction Methods

### v1

@@ 4 Main Results and Related Discussions
Second, Corollary 1 proves that more aggressive acceleration (i.e., larger $\alpha_{t}$) does not necessarily lead to better complexity in Algorithm 1 under the desirable constraint b=1b=1151515By desirable, we mean that smaller per-iteration cost is desirable if such choice does not deteriorate the complexity., while it is the trade-off between acceleration and variance reduction that actually matters. Since we take $C_{2}=1$, then $\Delta_{2}<1$ under the condition (4), which means that the largest allowable acceleration $\alpha_{t}=\mathcal{O}(t)$ (which corresponds to $\alpha=1$) can not provide best achievable complexity in either cases of (5) [^16] Note that $C_{2}=1$ is a desirable mild constant factor in the complexity. One can also easily check that $\Delta_{2}<1$ for a wide range of $\epsilon$ when $C_{2}>1$ takes a mild constant value..
@@
This phenomenon somehow contradicts with the conventional choice in acceleration methods. Typically, since the convergence rate of an acceleration method is $1/\alpha_{T}^{2}$ where $T$ is the total iteration number, larger $\alpha_{t}$ setting is favourable as long as it is allowable. For example, in stochastic optimization problems where the gradient oracle contains stochastic noise, large momentum parameters may fail since the complexity explodes or the convergence fails. It is in such cases that the momentum parameter is then reduced in order to adapt to the inexactness of the gradient estimator[42].
In contrast, all $\alpha\in[0,1]$ settings are allowable in Algorithm 1, but we actively choose to reduce the acceleration strength and successfully get compensated by the reduction of the per-iteration cost in expectation. As a result, Corollary 1 proves that such trade-off can be utilized to get near-optimal complexity in expectation, while adopting the most aggressive acceleration under high-cost-based variance reduction is theoretically inferior.
Third, when the target accuracy $\epsilon$ can be determined in advance and the condition (4) is satisfied, the parameter setting (5) has its own merits. Concretely, since the mini-batch size $b=1$ is the smallest possible choice and $p_{t}$ is smaller than that under larger $\alpha$ choice (e.g., $\alpha=1$), Algorithm 1 under (5) achieves the near-optimal complexity with less per-iteration cost, which is desirable in the finite-sum optimization.
We believe that the proposed active acceleration-variance reduction trade-off is of theoretical value, and may be of general interest beyond the considered problem setting and methods of this article.
@@
Finally, without requiring the condition (4) and the knowledge of $\epsilon$, Algorithm 1 achieves the near-optimal complexity in expectation, which significantly improves upon the previous best $\mathcal{O}(n+n/\sqrt{\epsilon})$ complexity of single-loop VR methods [30] under the same conditions, as the following proposition states.
@@
Under the setting of Theorem 2, setting $\alpha=1$ and $b=\mathcal{O}(\sqrt{n})$, the complexity in expectation of Algorithm 1 is $\widetilde{\mathcal{O}}(n+\sqrt{n}/\sqrt{\epsilon})$.
5 Conclusions
We propose the active acceleration-variance reduction trade-off in single-loop VR methods solving the composite general convex optimization problem. At the core of this idea is the proposed $p_{t}-\alpha_{t}$ correspondence, named the Harmonia technique, which essentially quantitatively determines how much variance reduction is needed to perform certain level of acceleration, or vise versa. The key parameter $\alpha\in[0,1]$ controls the acceleration-variance reduction trade-off and through which essentially controls the trade-off between convergence rate and per-iteration cost continuously in the proposed unifying algorithmic framework.
The continuous nature and flexibility of $\alpha\in[0,1]$ brings several desirable merits. First, several known complexity of some quite different methods are re-discovered under the proposed unifying framework Katyusha-H. Second, under an extra mild condition, Katyusha-H can achieve the near-optimal complexity $\widetilde{\mathcal{O}}(n+\sqrt{n}/\sqrt{\epsilon})$ for $\alpha$ belonging to certain explicitly given interval with the smallest possible mini-batch size $b=1$ and decreasing per-iteration cost in expectation (when $\alpha<1$). Third, setting $\alpha=1$ and $b=\mathcal{O}(\sqrt{n})$, Katyusha-H achieves the complexity $\widetilde{\mathcal{O}}(n+\sqrt{n}/\sqrt{\epsilon})$ under standard assumptions without knowledge of the target accuracy, which significantly improves upon previous best $\mathcal{O}(n+n/\sqrt{\epsilon})$ of single-loop VR methods and matches the theoretical lower bound $\Omega(n+\sqrt{n}/\sqrt{\epsilon})$ up to logarithmic factors.
We believe that the proposed idea and technique may be of general interest beyond the considered problem in this work.
Proof of Lemma 1. We begin by proving the following key inequality:
@@ 5 Conclusions
then inequality (7) is equivalent to $g(t)\geq 0$ for $t\geq 18$, which will be proved by a monotonicity argument. The simple case $\alpha=1$ will be discussed separately. We prove $g(17)\geq 0$ for $\alpha\in\left(0,1\right)$ as follows. In the $\alpha\in\left(0,1/2\right]$ case, note that
@@
where in the first inequality the concavity inequality $17^{\alpha}\geq(1/2)\left(16^{\alpha}+18^{\alpha}\right)$ is used, and the second inequality is by the fact that $r_{1}^{\alpha}-r_{2}^{\alpha}$ is decreasing with respect to (w.r.t.) $\alpha$, where $1<r_{1}<r_{2}$ (this can be shown by taking derivative w.r.t. $\alpha$). Simple calculation shows that $g(17)\geq 0$ in this case. Similarly, in the $\alpha\in\left(1/2,3/4\right]$ case,
@@
The following lemma mimics Lemma $2.4$ of [8].
@@
where in the second equality we applied (14), the third equality can be derived in a way similarly to (14) and we refer to Lemma $18$ and Lemma $19$ of [30], in the last inequality the following inequality is applied:
@@
where $g:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is a L-smooth convex function, and this inequality is a standard tool which can be found in, e.g., [43]. ∎
With Lemma 1 and Lemma 2, the convergence theorem can be proved as follows.
Proof of Theorem 1. Note that
@@
$+\alpha_{t}\langle\nabla f(x_{t+1}),x_{t+1}-z_{t}\rangle.$ (15)
where in the first inequality we applied the convexity of $f$, and the second equality is by the unbiasedness of the estimator $\tilde{g}_{t+1}$. The three terms in the RHS of (15) are separately bounded as follows. For the first term,
@@
$+\alpha_{t}\left(l(x^{*})-l(z_{t+1})\right).$ (17)
where $\bar{g}_{t+1}$ is a subgradient of $l$ at $z_{t+1}$, and in the first equality we use the equivalent update formula [^18] This is derived from the optimality of $z_{t+1}$ in the optimization problem $z_{t+1}=\arg\min_{z\in\mathbb{R}^{d}}\left\{\frac{1}{2\alpha_{t}\eta}\|z-z_{t}\|^{2}+\langle\tilde{g}_{t+1},z\rangle+l(z)\right\},$ which is a standard result and can be found in, e.g., [44]. of $z_{t+1}$, i.e., $z_{t+1}=z_{t}-\alpha_{t}\eta\left(\tilde{g}_{t+1}+\bar{g}_{t+1}\right)$, in the last inequality we use the convexity of $l$ and the three-point property $\langle a-b,b-c\rangle=(1/2)\cdot(\|a-c\|^{2}-\|a-b\|^{2}-\|b-c\|^{2}),\ \forall a,b,c\in\mathbb{R}^{d}$. Recall $x_{t+1}=\tau_{t}z_{t}+\xi w_{t}+(1-\xi-\tau_{t})y_{t}$ in Algorithm 1 and the setting $\tau_{t}\alpha_{t}=1$, we have
@@
Proof of Theorem 2. We divide the proof into cases where the $\{\alpha_{t}\}$ sequence has different growing rates, which is determined by $\alpha$ in Lemma 1. In all cases, we always investigate the convergence of the quantity $\mathbb{E}\left(F(w_{t})-F(x^{*})\right)$.
@@
where the last inequality is by $\xi\in(0,1)$ and $\sum_{t=2}^{\lceil 1/\epsilon\rceil}(1/t)\leq\int_{1}^{\lceil 1/\epsilon\rceil}(1/t)\text{d}t$. Hence in this case, the optimal choice of the mini-batch size is $b^{*}=1$, and the IFO complexity in expectation is $\mathcal{O}\left(n\log(1/\epsilon)+1/\epsilon\right)$.
@@
then for any $\alpha\in\left(0,1\right]$, when $t\geq 17$, we have $\tilde{\alpha}_{0}+\alpha_{0}^{2}-\alpha_{t}^{2}+\sum_{j=1}^{t}\alpha_{j}\geq\tilde{a}_{\alpha}t^{\alpha+1}$. Since $T_{\alpha}$ admits a uniform constant upper bound $17$ for any $\alpha\in\left(0,1\right]$, one can find a $\tilde{\tilde{a}}_{\alpha}>0$ (which is free from problem dependent parameters and possibly smaller than $\tilde{a}_{\alpha}$) so that $\tilde{\alpha}_{0}+\alpha_{0}^{2}-\alpha_{t}^{2}+\sum_{j=1}^{t}\alpha_{j}\geq\tilde{\tilde{a}}_{\alpha}t^{\alpha+1}$ for all $t\in\mathbb{N}^{+}$. Hence by Theorem 1, Algorithm 1 needs $T=\mathcal{O}\left(1/\epsilon^{1/\left(\alpha+1\right)}\right)$ iterations to reach an $\epsilon$-accurate solution of (1). To estimate $\mathbb{E}C_{\text{comp}}=\left(b/\epsilon^{1/\left(\alpha+1\right)}+n\sum_{t=1}^{\lceil 1/\epsilon^{1/\left(\alpha+1\right)}\rceil}p_{t}\right)$, the update probability $p_{t}$ is bounded as follows
@@
where $\xi=1/(bc)$ by our setting and $t\geq 17$. Note that although for $t<17$ the corresponding $p_{t}$ is not upper bounded carefully, we can always upper bound it by $1$, which at most leads to an addend being constant multiple of $n$ in the complexity, i,e., $\mathcal{O}(n)$. For $t\geq 17$, the inequality (26) implies
$p_{t}\leq\tilde{a}_{\alpha}^{-1}a_{\alpha}^{2}\xi\frac{1}{t^{1-\alpha}}+\tilde{a}_{\alpha}^{-1}a_{\alpha}\frac{1}{t}.$
@@
The quantity $Q_{1}(\epsilon,\alpha)$ is essentially upper bounded by $(1/\alpha)\cdot(1/\epsilon^{(\alpha/(\alpha+1))})$, which explodes when $\alpha\rightarrow 0^{+}$, hence the $\alpha\in\left(0,1/2\right]$ case is further handled as follows. Without loss of generality, assume $\epsilon\in(0,1)$, then for $\alpha\in\left(0,\min\left\{\hat{\alpha},1/10\right\}\right]$,
$Q_{1}(\epsilon,\alpha)=\sum_{t=17}^{\lceil 1/\epsilon^{1/(\alpha+1)}\rceil}\frac{t^{\alpha}}{t}\leq\sum_{t=17}^{\lceil 1/\epsilon\rceil}\frac{\left(\lceil 1/\epsilon\rceil\right)^{\hat{\alpha}}}{t}=\mathcal{O}\left(\log\left(1/\epsilon\right)\right).$
For $\alpha\in\left(\min\left\{\hat{\alpha},1/10\right\},1/2\right]$, the quantity $1/\alpha$ in the RHS of (27) can be upper bounded by $\max\left\{1/\hat{\alpha},10\right\}$, which is at the order of $\mathcal{O}(\log(1/\epsilon))$ [^19] For $\alpha\in\left(0,1/2\right]$ which is not very close to 0, e.g., $\alpha=1/4$, one can directly and safely apply (27).. Then the complexity in expectation is
$\mathbb{E}C_{\text{comp}}=\begin{cases}\mathcal{O}\left(\frac{b}{\epsilon^{1/(\alpha+1)}}+n\log(1/\epsilon)\right),\quad&\alpha\in\left(0,\min\left\{\hat{\alpha},1/10\right\}\right],\\ \mathcal{O}\left(\frac{b}{\epsilon^{1/(\alpha+1)}}+\frac{n}{b}\frac{1}{\alpha}\frac{1}{\epsilon^{\alpha/(\alpha+1)}}+n\log(1/\epsilon)\right),\quad&\alpha\in\left(\min\left\{\hat{\alpha},1/10\right\},1\right].\end{cases}$

### v2

@@ 4 Main Results and Related Discussions
Second, Corollary 1 proves that more aggressive acceleration (i.e., larger $\alpha_{t}$) does not necessarily lead to better complexity in Algorithm 1 under the mini-batch size constraint $b=1$, while it is the trade-off between acceleration and variance reduction that actually matters. Since we take $C_{2}=1$, then $\Delta_{2}<1$ under the condition (4), which means that the largest allowable acceleration $\alpha_{t}=\mathcal{O}(t)$ (which corresponds to $\alpha=1$) can not provide best achievable complexity in either cases of (5) [^9] Note that $C_{2}=1$ is a desirable mild constant factor in the complexity. One can also easily check that $\Delta_{2}<1$ for a wide range of $\epsilon$ when $C_{2}>1$ takes a mild constant value..
@@
This phenomenon somehow contradicts with the conventional choice in acceleration methods for solving noise-absent problems. Typically, since the convergence rate of an acceleration method is $1/\alpha_{T}^{2}$ where $T$ is the total iteration number, larger $\alpha_{t}$ setting is favourable as long as it is allowable.
In stochastic optimization problems where the gradient oracle contains stochastic noise, large momentum parameters may fail since the complexity explodes or the convergence fails. It is in such cases that the momentum parameter is then reduced in order to adapt to the inexactness of the gradient estimator [gupta2024nesterov]. For another example, when the gradient estimate satisfies the absolute error condition, the convergence bound of accelerated methods consists of the vanishing term and the accumulated error [liu2025nonasymptoticanalysisacceleratedmethods]. Since more aggressive acceleration decreases the vanishing term while increases the accumulated error, the acceleration strength is tuned to achieve the ideal trade-off in the finite-horizon setting where the iteration number is prefixed.
In contrast, in the considered noise-absent finite-sum problem, all $\alpha\in[0,1]$ settings are allowable in Algorithm 1 and there is no accumulated error term, but we actively choose to reduce the acceleration strength and successfully get compensated by the reduction of the per-iteration cost in expectation. The analysis shows that in Algorithm 1, acceleration comes with implicit price of heavier variance reduction (achieved by more per-iteration cost), hence the trade-off should be considered in certain scenarios even when the gradient oracle is noise-free.
@@
Finally, Algorithm 1 achieves the near-optimal complexity in expectation under standard assumptions, which significantly improves upon the previous best $\mathcal{O}(n+n/\sqrt{\epsilon})$ complexity of single-loop VR methods [driggs2022accelerating] and matches the lower bound $\Omega(n+\sqrt{n}/\sqrt{\epsilon})$ up to a logarithmic factor.
@@
Under the setting of Theorem 2, setting $\alpha=1$ and $b=\mathcal{O}(\sqrt{n})$, then
$\lim_{t\rightarrow\infty}\mathbb{E}\left[F(y_{t})-F(x^{*})\right]=0,\ \lim_{t\rightarrow\infty}\mathbb{E}\left[F(w_{t})-F(x^{*})\right]=0,$
@@
the complexity in expectation of Algorithm 1 (for both sequences $\{\mathbb{E}[F(y_{t})]\}_{t}$ and $\{\mathbb{E}[F(w_{t})]\}$) is $\widetilde{\mathcal{O}}(n+\sqrt{n}/\sqrt{\epsilon})$.
5 Numerical Experiments
The results of some numerical experiments are given to demonstrate the efficiency of the proposed method and validate other theoretical findings. We consider the following $l_{1}$-regularized logistic regression problem
$\mathop{\rm{min}}_{x\in\mathbb{R}^{d}}\quad F(x)=\frac{1}{n}\sum_{i=1}^{n}\log[1+\exp(-b_{i}a_{i}^{T}x)]+\frac{\lambda}{2}\|x\|_{1},$ (6)
where $a_{i}\in\mathbb{R}^{d}$ and $b_{i}\in\{-1,1\}$ are the feature vector and class label of the $i$th sample, respectively, and $\lambda>0$ is a weighting parameter. The adopted binary classification datasets ijcnn1, w8a, rcv1 are downloaded from the LIBSVM [^10] Online available at https://www.csie.ntu.edu.tw/~cjlin/libsvmtools/datasets/. database. All the tested algorithms are run $10$ times, and the best performance of these algorithms is reported.
5.1 Comparison with Other Methods
In this subsection, we compare Algorithm 1 with FISTA, Katyusha (for non-strongly convex problems) and VRADA for solving (6), where $\lambda=10^{-4}$. The Euclidean norm of each data vector in the datasets are normalized to be $1$, and hence $0.25$ is an upper bound of the smoothness parameter $L$ of the smooth part of (6).
For the four algorithms we compare, the common parameter to tune is the parameter with respect to (w.r.t.) the Lipschitz constant, which is tuned in $\{0.0125,0.025,0.05,0.1,0.25,0.5\}$. The inner loop length $m$ of Katyusha and VRADA are set as $m=2n$, as suggested by [allen2018katyusha, song2020variance]. For Algorithm 1, we set $\alpha=1$ and the mini-batch size $b$ as $b=\lceil 1/\sqrt{n}\rceil$, as required in Corollary 2. Other parameters of Algorithm 1 are set strictly according to the theoretical requirements, where $\tilde{\alpha}_{0}$ is set as $\tilde{\alpha}_{0}=\max\{\xi\alpha_{1}^{2},36\}$ which satisfies the theoretical requirement in Lemma 1 and prevents frequent computation of full gradient at the beginning of the optimization process.
The comparison results of tested algorithms are shown in Fig. 1. In Fig. 1, the $x$-axis denotes the number of full gradient evaluations, the $y$-axis denotes the objective function value at the checkpoints of the VR methods (and the iterates of the FISTA method).
It can be seen from Fig. 1 that, Algorithm 1 converges to a moderate accuracy faster than other tested algorithms, saving a substantial amount of full gradient evaluations compared to FISTA, Katyusha, and VRADA, in the sense that the ratio between required epochs of Algorithm 1 and that of other methods is small. This feature is favourable in certain scenarios, e.g., in large-scale machine learning tasks where a solution of moderate accuracy is required.
(a) ijcnn1
(b) a9a
(c) w8a
Figure 1: Performance comparison for solving regularized logistic regression among different algorithms
5.2 Empirical Investigation of the Acceleration-Variance Reduction Trade-Off
In this subsection, we investigate the proposed acceleration-variance reduction trade-off numerically. Algorithm 1 with different $\alpha$ settings are used for solving (6), where $\lambda=10^{-4}$. The Euclidean norm of each data vector in the datasets are normalized to be $1$, and hence $0.25$ is an upper bound of the smoothness parameter $L$ of the smooth part of (6).
For Algorithm 1 with different $\alpha$ settings, we use the Lipschitz constant estimate $0.25$ and set the mini-batch size $b$ as $b=1$. Other parameters are set the same and strictly satisfy the theoretical requirements of Corollary 1 to highlight and investigate the effect of the key parameter $\alpha$.
The comparison results are shown in Fig. 2. In Fig. 2, the $x$-axis denotes the number of full gradient evaluations, the $y$-axis denotes the gap between the objective function value at the checkpoint and the approximate minimum of the objective function value. The intersections on the $x$-axis indicate required full gradient evaluations to reach an $\epsilon$-accurate solution, where $\epsilon$ is the prefixed target accuracy.
In Fig. (2(a)), (2(b)), and (2(c)), the target accuracy is $\epsilon=1/\sqrt{n}$ where $n$ is the number of samples of the dataset. In this low-accuracy case, the optimal $\alpha$ value suggested in Corollary 1 is $\alpha=0$. In Fig. (2(d)), (2(e)), and (2(f)), the target accuracy is $\epsilon=1/(5n)$. In this case, the $\alpha$ is set as
$\alpha=\alpha^{*}=\frac{1-\frac{\log(n)}{\log(1/\epsilon)}}{1+\frac{\log(n)}{\log(1/\epsilon)}},$
which satisfies the theoretical requirement of Corollary 1.
It can be seen from Fig. 2 that, Algorithm 1 with theoretically suggested (small) $\alpha$ setting significantly outperforms Algorithm 1 with large $\alpha$ setting. Algorithm 1 with large $\alpha$ setting, i.e., aggressive acceleration strength, requires relatively frequent full gradient computation. This heavier variance reduction, although allows more aggressive acceleration, ultimately leads to inferior empirical performance for the given accuracy, which validates our theoretical results in Corollary 1. It also follows that it is beneficial to consider the proposed acceleration-variance reduction trade-off rather than acceleration alone in certain scenarios.
(a) ijcnn1
(b) a9a
(c) w8a
(d) ijcnn1
(e) a9a
(f) w8a
Figure 2: Performance comparison of Algorithm 11 under different $\alpha$ settings
6 Conclusion
The trade-off between convergence rate and per-iteration cost, which essentially determine the complexity of an algorithm, is important in (structured) finite-sum optimization. Handling such trade-off typically involves the usage of different algorithmic structure, iteration direction, and other optimization technique. In this work, we propose a unified single-loop variance-reduced acceleration framework. By relating checkpoint update probabilities to momentum parameters and alternating the growth rate of the momentum parameters, the proposed simple framework exhibits a continuous trade-off between convergence rate and per-iteration cost, obtaining a series of new complexity and recovering some known complexity of distinct existing methods as special cases.
The proposed framework allows us to explore the trade-off between acceleration and variance reduction, which is key to achieve the aforementioned rate-cost trade-off and is itself a novel topic in noise-absent finite-sum optimization. The analysis shows that in the proposed framework, acceleration comes with implicit price of heavier variance reduction (achieved by increasing per-iteration cost), which renders the largest allowable acceleration strength not necessarily optimal and highlights the introduced acceleration-variance reduction trade-off in certain scenarios.
The proposed framework also leads to complexity improvement in the family of single-loop variance reduction methods. Under standard assumptions, our method achieves the near-optimal complexity $\tilde{\mathcal{O}}(n+\sqrt{n}/\sqrt{\epsilon})$, which significantly improves upon previous best complexity $\mathcal{O}(n+n/\sqrt{\epsilon})$ in this family of methods for solving the composite general convex problem.
Appendix A Proof of Lemma 1
We begin by proving the following key inequality:
@@ Appendix A Proof of Lemma 1
then inequality (8) is equivalent to $g(t)\geq 0$ for $t\geq 18$, which will be proved by monotonicity argument. The simple case $\alpha=1$ will be discussed separately. We prove $g(17)\geq 0$ for $\alpha\in\left(0,1\right)$ as follows. In the $\alpha\in\left(0,1/2\right]$ case, note that
@@
where in the first inequality the concavity inequality $17^{\alpha}\geq(1/2)\left(16^{\alpha}+18^{\alpha}\right)$ is used, and the second inequality is by the fact that $r_{1}^{\alpha}-r_{2}^{\alpha}$ is decreasing w.r.t. $\alpha$, where $1<r_{1}<r_{2}$ (this can be shown by taking derivative w.r.t. $\alpha$). Simple calculation shows that $g(17)\geq 0$ in this case. Similarly, in the $\alpha\in\left(1/2,3/4\right]$ case,
@@
Appendix B Bounding the Variance of the SVRG Estimator
The following lemma mimics Lemma $2.4$ of [allen2018katyusha].
@@ Appendix B Bounding the Variance of the SVRG Estimator
where in the second equality we apply (15), the third equality can be derived in a way similarly to (15) and we refer to Lemma $18$ and Lemma $19$ of [driggs2022accelerating], in the last inequality the following property is applied:
@@
where $g:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is a L-smooth convex function, and this property is a standard tool which can be found in [nesterov2018lectures]. ∎
Appendix C Proof of Theorem 1
Note that
@@ Appendix C Proof of Theorem 1
$+\alpha_{t}\langle\nabla f(x_{t+1}),x_{t+1}-z_{t}\rangle,$ (16)
where in the first inequality we apply the convexity of $f$, and the second equality is by the unbiasedness of the estimator $\tilde{g}_{t+1}$. The three terms in the RHS of (16) are separately bounded as follows. For the first term,
@@
$+\alpha_{t}\left(l(x^{*})-l(z_{t+1})\right),$ (18)
where $\bar{g}_{t+1}$ is a subgradient of $l$ at $z_{t+1}$, and in the first equality we use the equivalent update formula [^12] This is derived from the optimality of $z_{t+1}$ in the optimization problem $z_{t+1}=\arg\min_{z\in\mathbb{R}^{d}}\left\{\frac{1}{2\alpha_{t}\eta}\|z-z_{t}\|^{2}+\langle\tilde{g}_{t+1},z\rangle+l(z)\right\},$ which is a standard result and can be found in [doi:10.1137/1.9781611974997]. of $z_{t+1}$, i.e., $z_{t+1}=z_{t}-\alpha_{t}\eta\left(\tilde{g}_{t+1}+\bar{g}_{t+1}\right)$, in the last inequality we use the convexity of $l$ and the three-point property $\langle a-b,b-c\rangle=(1/2)\cdot(\|a-c\|^{2}-\|a-b\|^{2}-\|b-c\|^{2}),\ \forall a,b,c\in\mathbb{R}^{d}$. Recall $x_{t+1}=\tau_{t}z_{t}+\xi w_{t}+(1-\xi-\tau_{t})y_{t}$ in Algorithm 1 and the setting $\tau_{t}\alpha_{t}=1$, we have
@@
Appendix D Proof of Theorem 2
We divide the proof into cases where the $\{\alpha_{t}\}$ sequence has different growing rate, which is determined by $\alpha$ in Lemma 1. In all cases, we always consider the convergence of the quantity $\mathbb{E}\left(F(w_{t})-F(x^{*})\right)$.
@@ Appendix D Proof of Theorem 2
where the last inequality is by $\xi\in(0,1)$ and $\sum_{t=2}^{\lceil 1/\epsilon\rceil}(1/t)\leq\int_{1}^{\lceil 1/\epsilon\rceil}(1/t)\text{d}t$. In addition, when $n<\lceil 1/\epsilon\rceil$,
$\frac{b}{\epsilon}+n\sum_{t=1}^{\lceil 1/\epsilon\rceil}p_{t}$ $\leq\frac{b}{\epsilon}+\frac{n(36\xi+6)}{6}\sum_{t=1}^{n}\frac{1}{t}+\frac{n(36\xi+6)}{6}\sum_{t=n+1}^{\lceil 1/\epsilon\rceil}\frac{1}{t}$
$\leq\frac{b}{\epsilon}+7n+7n\log(n)+7n\sum_{t=n+1}^{\lceil 1/\epsilon\rceil}\frac{1}{n}$
$\leq\frac{b}{\epsilon}+7n\log(n)+7\lceil 1/\epsilon\rceil.$
In conclusion, in the $\alpha=0$ case, the optimal choice of the mini-batch size is $b^{*}=1$, and the IFO complexity in expectation is $\mathcal{O}\left(n\min\{\log(1/\epsilon),\log(n)\}+1/\epsilon\right)$.
@@
then for any $\alpha\in\left(0,1\right]$, when $t\geq 17$, we have $\tilde{\alpha}_{0}+\alpha_{0}^{2}-\alpha_{t}^{2}+\sum_{j=1}^{t}\alpha_{j}\geq\tilde{a}_{\alpha}t^{\alpha+1}$. Since $T_{\alpha}$ admits a uniform constant upper bound $17$ for any $\alpha\in\left(0,1\right]$, one can find a $\tilde{\tilde{a}}_{\alpha}>0$ (which is free from problem dependent parameters and possibly smaller than $\tilde{a}_{\alpha}$) so that $\tilde{\alpha}_{0}+\alpha_{0}^{2}-\alpha_{t}^{2}+\sum_{j=1}^{t}\alpha_{j}\geq\tilde{\tilde{a}}_{\alpha}t^{\alpha+1}$ for all $t\in\mathbb{N}^{+}$. Hence by Theorem 1, Algorithm 1 needs $T=\mathcal{O}\left(1/\epsilon^{1/\left(\alpha+1\right)}\right)$ iterations to reach an $\epsilon$-accurate solution of (1). To estimate
$\mathbb{E}C_{\text{comp}}=b/\epsilon^{1/\left(\alpha+1\right)}+n\sum_{t=1}^{\lceil 1/\epsilon^{1/\left(\alpha+1\right)}\rceil}p_{t},$ (27)
the update probability $p_{t}$ is bounded as follows
@@
where $t\geq 17$ and $\xi=1/(bc)$ by our setting. Note that although for $t<17$ the corresponding $p_{t}$ is not upper bounded carefully, we can always upper bound it by $1$, which at most leads to an addend being constant multiple of $n$ in the complexity, i,e., $\mathcal{O}(n)$. For $t\geq 17$, the inequality (27) implies
$p_{t}\leq\tilde{a}_{\alpha}^{-1}a_{\alpha}^{2}\xi\frac{1}{t^{1-\alpha}}+\tilde{a}_{\alpha}^{-1}a_{\alpha}\frac{1}{t}.$ (29)
@@
The quantity $Q_{1}(\epsilon,\alpha)$ is essentially upper bounded by $(1/\alpha)\cdot(1/\epsilon^{(\alpha/(\alpha+1))})$, which explodes when $\alpha\rightarrow 0^{+}$, hence the $\alpha\in\left(0,1/2\right]$ case is further handled as follows. Without loss of generality, assume $\epsilon\in(0,1)$, then for $\alpha\in\left(0,\min\left\{\hat{\alpha},1/10\right\}\right]$, when $\hat{\alpha}=\log(2)/\log(\lceil 1/\epsilon\rceil)$, i.e., $n>\lceil 1/\epsilon\rceil$,
$Q_{1}(\epsilon,\alpha)=\sum_{t=17}^{\lceil 1/\epsilon^{1/(\alpha+1)}\rceil}\frac{t^{\alpha}}{t}\leq\sum_{t=17}^{\lceil 1/\epsilon^{(1/(\alpha+1))}\rceil}\frac{\left(\lceil 1/\epsilon\rceil\right)^{\hat{\alpha}}}{t}=\mathcal{O}(\log(1/\epsilon)).$
When $\hat{\alpha}=\log(2)/\log(n)$, i.e., $n<\lceil 1/\epsilon\rceil$, if $\lceil 1/\epsilon^{1/(\alpha+1)}\rceil\leq n<\lceil 1/\epsilon\rceil$,
$Q_{1}(\epsilon,\alpha)\leq\sum_{t=17}^{n}\frac{t^{\alpha}}{t}\leq\sum_{t=17}^{n}\frac{n^{\hat{\alpha}}}{t}=\mathcal{O}(\log(n)),$
if $n<\lceil 1/\epsilon^{1/(\alpha+1)}\rceil$,
$Q_{1}(\epsilon,\alpha)$ $\leq\sum_{t=17}^{\lceil 1/\epsilon^{1/(\alpha+1)}\rceil}\frac{1}{t^{1-\alpha}}=\sum_{t=17}^{n}\frac{1}{t^{1-\alpha}}+\sum_{t=n+1}^{\lceil 1/\epsilon^{1/(\alpha+1)}\rceil}\frac{1}{n^{1-\alpha}}$
$\leq\sum_{t=17}^{n}\frac{n^{\alpha}}{t}+\sum_{t=n+1}^{\lceil 1/\epsilon^{1/(\alpha+1)}\rceil}\frac{n^{\alpha}}{n}\leq\sum_{t=17}^{n}\frac{n^{\hat{\alpha}}}{t}+\sum_{t=n+1}^{\lceil 1/\epsilon^{1/(\alpha+1)}\rceil}\frac{n^{\hat{\alpha}}}{n}$
$=\mathcal{O}\left(\log(n)+\frac{1}{n}\frac{1}{\epsilon^{1/(\alpha+1)}}\right).$
Hence for $\alpha\in\left(0,\min\left\{\hat{\alpha},1/10\right\}\right]$,
$Q_{1}(\epsilon,\alpha)=\mathcal{O}\left(\left(1+\frac{1}{n}\frac{1}{\epsilon^{1/(\alpha+1)}}\right)\min\left\{\log(n),\log(1/\epsilon)\right\}\right).$ (32)
For $\alpha\in\left(\min\{\hat{\alpha},1/10\},1\right]$, we have $1/\alpha=\mathcal{O}(\min\{\log(n),\log(1/\epsilon)\})$, then by (30),
$Q_{1}(\epsilon,\alpha)=\mathcal{O}\left(\frac{1}{\epsilon^{\alpha/(\alpha+1)}}\min\left\{\log(n),\log(1/\epsilon)\right\}\right).$ (33)
By using the argument in the $\alpha=0$ case, it can be proved that when $n<\lceil 1/\epsilon^{1/(\alpha+1)}\rceil$, $Q_{2}(\epsilon,\alpha)=\mathcal{O}(\log(n)+(1/n)\epsilon^{(-1)/(\alpha+1)})$. Hence for all $\alpha\in\left(0,1\right]$,
$Q_{2}(\epsilon,\alpha)=\mathcal{O}\left(\left(1+\frac{1}{n}\frac{1}{\epsilon^{1/(\alpha+1)}}\right)\min\left\{\log(n),\log(1/\epsilon)\right\}\right).$ (34)
Combining (27), (29), (32), (33), and (34), and noting that $\xi=1/(bc)$, the complexity in expectation of Algorithm 1 is
$\mathbb{E}C_{\text{comp}}=\begin{cases}\mathcal{O}\left(\frac{b}{\epsilon^{1/(\alpha+1)}}+n\min\left\{\log(n),\log(1/\epsilon)\right\}\right),\quad&\alpha\in\left(0,\min\left\{\hat{\alpha},1/10\right\}\right],\\ \mathcal{O}\left(\frac{b}{\epsilon^{1/(\alpha+1)}}+\left(\frac{n}{b}\frac{1}{\epsilon^{\alpha/(\alpha+1)}}+n\right)\min\left\{\log(n),\log(1/\epsilon)\right\}\right),\quad&\alpha\in\left(\min\left\{\hat{\alpha},1/10\right\},1\right].\end{cases}$